OpenStudy (anonymous):

Simplify the following (Will write it in equation below). Can you explain the steps in solving it?

OpenStudy (anonymous):

\[\frac{ -1 +3i }{ 4-i }\]

OpenStudy (anonymous):

Let z = 4 - i, and find the complement:  Multiply your fraction by zbar/zbar, since zbar/zbar = 1:  When you expand this out, you'll end up with a complex number on top and a real number as your denominator. It multiplies out like any normal pair of brackets would (using FOIL) - when you get -(i^2), it will become -(sqrt(-1))^2 = +1 Let me know if you want help with the expanding part.